OpenStudy (anonymous):

Multiply or divide as indicated. Leave your answer with no factors in the denominator. x^-7 /x^-8

OpenStudy (anonymous):

x

OpenStudy (anonymous):

x^-7 /x^-8=(x^-7 )(x^+8)=x^1=x ans..